OpenStudy (anonymous):

When you multiply a function by -1, what is the effect on its graph? A. The graph flips over the x-axis. B. The graph flips over the line y = x. C. The graph flips over the y-axis.

OpenStudy (whpalmer4):

Why don't you plot the function \[y = x^2\]from \(x=-2\) to \(x=2\) Next, plot the function \[y = -1*x^2\]over the same range. Discuss.